# Editing Factual Knowledge and Explanatory Ability of Medical Large Language Models

### v2

@@ Abstract.
Model editing aims to precisely alter the behaviors of large language models (LLMs) in relation to specific knowledge, while leaving unrelated knowledge intact. This approach has proven effective in addressing issues of hallucination and outdated information in LLMs. However, the potential of using model editing to modify knowledge in the medical field remains largely unexplored, even though resolving hallucination is a pressing need in this area. Our observations indicate that current methods face significant challenges in dealing with specialized and complex knowledge in medical domain. Therefore, we propose MedLaSA, a novel Layer-wise Scalable Adapter strategy for medical model editing. MedLaSA harnesses the strengths of both adding extra parameters and locate-then-edit methods for medical model editing. We utilize causal tracing to identify the association of knowledge in neurons across different layers, and generate a corresponding scale set from the association value for each piece of knowledge. Subsequently, we incorporate scalable adapters into the dense layers of LLMs. These adapters are assigned scaling values based on the corresponding specific knowledge, which allows for the adjustment of the adapter’s weight and rank. The more similar the content, the more consistent the scale between them. This ensures precise editing of semantically identical knowledge while avoiding impact on unrelated knowledge. To evaluate the editing impact on the behaviours of LLMs, we propose two model editing studies for medical domain: (1) editing factual knowledge for medical specialization and (2) editing the explanatory ability for complex knowledge. We build two novel medical benchmarking datasets and introduce a series of challenging and comprehensive metrics. Extensive experiments on medical LLMs demonstrate the editing efficiency of MedLaSA, without affecting unrelated knowledge.  [^1] Our code and data will be publicly available.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: of the 33rd ACM International Conference on Information and Knowledge Management; October 21 – 25, 2024; Boise, Idaho, USAisbn: 978-1-4503-XXXX-X/18/06ccs: Applied computing Life and medical sciences
@@ 1. Introduction
The emergence of large language models (LLMs) has greatly promoted the development of natural language processing due to their powerful text generation capabilities (Zhao et al., 2023; Xu et al., 2023). Meanwhile, LLMs also garner increasing attention from researchers in healthcare field (Thirunavukarasu et al., 2023; He et al., 2023). Although LLMs have proven to be valuable tools, they may still provide outdated factual information or even experience hallucinations (Zhang et al., 2023; Li et al., 2023a; Ye et al., 2023), which is particularly concerning when deployed in the real-world medical scenarios (Hartvigsen et al., 2022; Feng et al., 2023). To solve this problem, considering the substantial cost of retraining an LLM, there has been an increasing interest in model editing (also known as knowledge editing) (Wang et al., 2023a). It seeks to modify the behaviors of LLMs by precisely manipulating a part of knowledge while ensuring unrelated knowledge unaffected (not the target of editing) (Zhang et al., 2024; Meng et al., 2022a, b; Wang et al., 2023a). Researches have demonstrated that the LLMs can serve as a knowledge base for storing factual information about the world (Petroni et al., 2019; Geva et al., 2022) and the knowledge in LLMs can be edited by understanding mechanization of knowledge retention within the transformer structure (Dai et al., 2022; Geva et al., 2021; Lv et al., 2024).
@@
We observed that these methods struggle to overcome challenges of specialized and complex knowledge in the medical domain (Karabacak and Margetis, 2023). As shown in Figure 1, LLMs may exhibit hallucinations and provide answers that do not align with the real world. It lacks specialized knowledge regarding the diseases that can be treated with amoxicillin (Chang et al., 2024; Kung et al., 2023), leading to incorrect responses such as ”stinging”. When dealing with complex knowledge, it is common to employ advanced prompting techniques (e.g., Chain-of-Thought (Lyu et al., 2023)), to facilitate logical reasoning and deduce rational conclusions. However, LLMs may fail to explain the rationale behind the impact of amoxicillin in treating tonsillitis, resulting in an erroneous reasoning chain. Current model editing methods either overlook the storage of complex medical knowledge across different layers in LLMs (Zheng et al., 2023) or introduce substantial modifications to original parameters, which consequently affect the model’s unrelated knowledge and lead to sub-optimal performance (Zhang et al., 2023). This research gap presents a question: Can model editing techniques effectively address challenges when integrating LLMs in the medical domain?
@@
Extensive experiments conducted on two comprehensive benchmarks demonstrate the superior performance of MedLaSA across a range of metrics, showing that it can achieve a higher editing success rate in specialized and complex medical knowledge contexts and avoid impacting irrelevant contexts. Our contributions are summarized as follows:
  • To the best of our knowledge, we are the first to propose model editing for enhancing factual knowledge and explanatory abilities in medical LLMs. And we construct two benchmarks with comprehensive evaluation metrics.
  • We introduce MedLaSA, a novel approach that dynamically adjusts the adapters across different layers of LLMs, which effectively addresses the challenges of model editing in specialized and complex medical knowledge.
  • We conduct extensive experimental analysis of specialized and complex medical knowledge editing, demonstrating that MedLaSA significantly outperforms the existing methods.
@@ 2. Related Work
We present current model editing works in two categories following Yao et al. (2023).
Memories or Additional Parameters.
The methods of with memories or additional parameters typically involve creating explicit memories to store the required knowledge for editing, or adding additional trainable parameters to LLMs for learning new knowledge (Yu et al., 2023; Dong et al., 2022; Hartvigsen et al., 2022). SERAC (Mitchell et al., 2022b) utilized explicit memory for storing edits and incorporated a scope classifier to understand the editing scope. Given a sample within the editing scope, it utilized a separate model to make edits, ensuring that the original model remains unaffected. Inspired by the in-context learning ability of LLMs, IKE (Zheng et al., 2023) designed demonstration formatting and organization strategies, including the copy, update, and retain templates, and retrieved relevant knowledge facts from the editing memories as demonstration inputs to guide the editing process. T-Patcher (Huang et al., 2023) retained all original parameters to preserve overall performance while adding trainable neuron patches to the last Feed-Forward Network (FFN) layer of a Transformer for handling sequential model editing. Despite their success, the above methods lack the exploration of the mechanics of knowledge storage in LLMs, which ultimately leads to poor performance in handling complex medical knowledge.
Modifying LLMs’ Parameters.
@@
The methods of this category aim to comprehend how knowledge is stored in LLMs and how it can be effectively altered by changing the parameters (De Cao et al., 2021; Geva et al., 2021; Wu et al., 2023). KN (Dai et al., 2022) proposed a knowledge attribution method to identify the neurons associated with specific knowledge without fine-tuning, updating facts, and erasing relations by directly modifying the corresponding parameters in FFN. MEND (Mitchell et al., 2022a) introduced auxiliary hyper-networks to transform the gradient during the fine-tuning process, and trained the hyper-networks to ensure edit success and locality when updating LLMs’ parameters. ROME (Meng et al., 2022a) applied Causal mediation analysis (Pearl, 2022; Vig et al., 2020) to identify decisive neuron activation and modify FFN weights by solving a least squares problem with a linear equality constraint using the Lagrange multiplier. As an extension of ROME (Meng et al., 2022a), MEMIT (Meng et al., 2022b) introduced a multi-layer algorithm to update multiple cases simultaneously. PMET (Li et al., 2023b) further improved MEMIT (Meng et al., 2022b) by simultaneously optimizing hidden states of self-attention and FFN. Despite impressive progress made by these methods, they often introduce significant modifications to the original parameters. Consequently, unrelated knowledge is affected, resulting in a noticeable impact on Locality and Fluency, as demonstrated in Section Experiments 4.
@@
Overall, both of them have their advantages and disadvantages. Neuron localization can achieve interpretability between parameters and knowledge, but sometimes optimizing too much can have negative impacts on unrelated knowledge (Meng et al., 2022a; Geva et al., 2021). On the other hand, adding extra parameters has a smaller impact on unrelated knowledge but lacks exploration of the mechanics of knowledge storage (Huang et al., 2023). Our proposed method, MedLaSA, aims to combine the strengths of both approaches. It identify the association of knowledge in neurons of different layers, and introduce scalable adapters into the dense layers, enables precise editing without affecting unrelated knowledge.
@@ 3.1. Prelimimaries
Figure 2. The overview of our proposed MedLaSA. We demonstrate the process of inputting editing, rephrased, and unrelated knowledge. The activation scale of adapters in different layers varies depending on the knowledge in the input question, ensuring that unrelated knowledge is not significantly impacted after the editing process.
@@ 3.3. MedLaSA
Our motivation lies in that different knowledge will have different scale sets. Similar contents typically share more consistent scales between them. As a result, it ensures accurate editing of identical knowledge while avoiding impact on unrelated knowledge. For instance, as illustrated in Figure 2, when data that needs editing is inputted, we increase its scale on the $i$-th layer and decrease its impact on the $j$-th layer, ensuring that knowledge updates occur on the $i$-th layer. Conversely, when unrelated data is input, the influence of the adapter on the $j$-th layer is activated and reduces the impact of the adapter on the $i$-th layer. Inspired by LoRA (Hu et al., 2021), we incorporate a parallel low-rank adapter into dense layers to enable parameter-efficient fine-tuning. Specifically, a pre-trained weight matrix $\boldsymbol{W}_{0}\in\mathbb{R}^{d\times k}$ from attention or MLP module is updated with two trainable matrics $\boldsymbol{B}\in\mathbb{R}^{d\times r}$ and $\boldsymbol{A}\in\mathbb{R}^{r\times k}$. For layer $l$, we have:
@@
The final rank of layer $l$ can be computed by multiplying hyper-parameter $r_{o}$: $r^{l}=\lceil r_{o}\times I^{l}_{r}\rceil$.
By computing scale values $\alpha$ and $r$, we can create a set of layer-wise scaling factors $\{r^{l},\alpha^{l}|l\in[1,L]\}$ for editing, rephrased, and unrelated knowledge, as shown in Figure 2. Besides, a Transformer block typically includes self-attention and MLP modules, whose effects can be separately analyzed with causal tracing. For example, to measure the impact matrix $\boldsymbol{M}_{attn}$ of the attention module, the MLP calculation is cut off and frozen in its corrupted run state, so that it is not affected by the insertion of a clean state. We conducted a thorough analysis to examine the influence of various weights and hyper-parameters on model editing, in Section 4.
@@ 3.4. Medical Model Editing Benchmarks
Explanation / Edit Target: Adiponectin is an abundant adipose-derived protein and enhances insulin sensitivity and lipid oxidation. Its levels are reduced in obesity Obesity is associated with significant disturbances in endocrine function. Hyper insulinemia and insulin resistance are the best known changes in obesity. Thyroxine, GH, and adiponectin have lipolytic effects, hence their levels are reduced in obesity.
@@ 3.4.1. Efficacy and Generality Data Construction
To build the MedCF dataset, we utilize a medical knowledge graph called DRKG (Ioannidis et al., 2020). For each triple $(h,r,t)$, we convert $h$ and $r$ into a question and $t$ into an answer, thus creating a QA pair. To evaluate the ability to edit knowledge with unknown prediction results of LLMs, we introduce counterfactual data for model training following ROME (Meng et al., 2022a). We randomly sample a negative entity $t_{*}$ from the knowledge graph and use it as a new answer, attempting to edit the original answer accordingly. For example in Table 1, the original triple “(Primaquine, side effect, Nausea)” is converted to the question: “What side effect is caused by Primaquine?” and the answer “Nausea”. But the editing target is indeed a counterfactual result “Stinging”.
@@
The data of Efficacy for MedCF and Generality for both MedCF and MedFE are generated by querying ChatGPT. The template for querying is shown in Table 2. This prompt template ensures that the generated content adheres to medical terminology while preserving the original meaning of the question. Different from MedCF, all the facts require paraphrasing for the MedFE dataset. We have examined all 17 types of relations and made efforts to ensure the rephrased sentences do not change from the original meaning.
@@
Table 2. Templates for querying ChatGPT to generate the Efficacy and Generality data. $h$, $r$, and $t$ denote the head entity, relation, and tail entity, respectively.
Figure 3. Knowledge type distribution of MedCF dataset.
@@
The QA pair of Locality ($x_{i}$,$y_{i}$) aims to assess the impact of model editing on unrelated knowledge. Previous benchmarks either employed out-of-distribution data (e.g., zsRE (Mitchell et al., 2022a)) or solely relied on data with the same ground truth (e.g., CounterFact (Meng et al., 2022a)). Nevertheless, we argue that a comprehensive evaluation of Locality is necessary to prevent the inadvertent modification of irrelevant knowledge and ensure high reliability of the medical domain. As we explained in Section 1, the post-edited model should be evaluated based on the following challenges: Target Distribution (TD), Entity Mapping (EM), Structural Similarity (SS), Textual Similarity (TS), and Consistent Topic (CT). To fulfill the requirements, we gathered relevant QA pairs for evaluation. The original triple of Efficacy QA pair is denoted as $(h,r,t)$. We employ the graph structure of KG for searching and building Locality sub-metrics. For Locality-TD, we sample new triples $(h_{td},r_{td},t)$ by fixing $t$ to investigate the shift of the ground-truth token. For Locality-EM, we sample new triples $(h,r_{e}m,t_{e}m)$ by fixing $h$ to create mappings between head and tail entities. For Locality-SS, we utilize the knowledge graph embedding method, RotatE (Sun et al., 2018), to learn the entity and relation embeddings from the graph structure. For Locality-TD, we use BioBERT (Lee et al., 2020) to convert $(h,r,t)$ triples into text embeddings, and then obtain similar triples $(h_{s},r_{s},t_{s})$ based on Structural Similarity and Textual Similarity, respectively. For Locality-CT, we utilize the source data from MedMCQA, which includes topic tags for each question, annotated by experts. These new triples are also converted into QA pairs, which enable a thorough analysis of the impacts of model editing techniques on unrelated knowledge in medical domain. It should be noted that the calculation for all sub-metrics of Locality is the same, as listed in Section 3.1.
We provide knowledge type distribution of MedCF and MedFE in Figure 3 and 4, respectively. The knowledge types in MedCF are categorized based on their relations, while the types in MedFE are categorized according to medical subjects. Data splits of MedCF and MedFE are shown in Table 3. We have attempted to ensure that different categories have similar proportions, but due to data limitations, some types have not been collected as extensively as others. More details about datasets can be found in our repository.

### v3

@@ Abstract.
Model editing aims to precisely alter the behaviors of large language models (LLMs) in relation to specific knowledge, while leaving unrelated knowledge intact. This approach has proven effective in addressing issues of hallucination and outdated information in LLMs. However, the potential of using model editing to modify knowledge in the medical field remains largely unexplored, even though resolving hallucination is a pressing need in this area. Our observations indicate that current methods face significant challenges in dealing with specialized and complex knowledge in medical domain. Therefore, we propose MedLaSA, a novel Layer-wise Scalable Adapter strategy for medical model editing. MedLaSA harnesses the strengths of both adding extra parameters and locate-then-edit methods for medical model editing. We utilize causal tracing to identify the association of knowledge in neurons across different layers, and generate a corresponding scale set from the association value for each piece of knowledge. Subsequently, we incorporate scalable adapters into the dense layers of LLMs. These adapters are assigned scaling values based on the corresponding specific knowledge, which allows for the adjustment of the adapter’s weight and rank. The more similar the content, the more consistent the scale between them. This ensures precise editing of semantically identical knowledge while avoiding impact on unrelated knowledge. To evaluate the editing impact on the behaviours of LLMs, we propose two model editing studies for medical domain: (1) editing factual knowledge for medical specialization and (2) editing the explanatory ability for complex knowledge. We build two novel medical benchmarking datasets and introduce a series of challenging and comprehensive metrics. Extensive experiments on medical LLMs demonstrate the editing efficiency of MedLaSA, without affecting unrelated knowledge.
@@
† Corresponding authors
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3679673isbn: 979-8-4007-0436-9/24/10ccs: Applied computing Life and medical sciences
@@ 1. Introduction
The emergence of large language models (LLMs) has greatly promoted the development of natural language processing due to their powerful text generation capabilities (Zhao et al., 2023; Li et al., 2024; Chen et al., 2024; Xu et al., 2023a). Meanwhile, LLMs also garner increasing attention from researchers in the healthcare field (Thirunavukarasu et al., 2023; He et al., 2023). Although LLMs have proven to be valuable tools, they may still provide outdated factual information or even experience hallucinations (Zhang et al., 2023; Li et al., 2023; Ye et al., 2023), which is particularly concerning when deployed in real-world medical scenarios (Hartvigsen et al., 2022; Feng et al., 2023; Liu et al., 2024). To solve this problem, considering the substantial cost of retraining an LLM, there has been an increasing interest in model editing (also known as knowledge editing) (Wang et al., 2023a). It seeks to modify the behaviors of LLMs by precisely manipulating a part of knowledge while ensuring unrelated knowledge is unaffected (not the target of editing) (Zhang et al., 2024; Meng et al., 2022a, b; Wang et al., 2023a). Researches have demonstrated that the LLMs can serve as a knowledge base for storing factual information about the world (Petroni et al., 2019; Tang et al., 2023; Geva et al., 2022) and the knowledge in LLMs can be edited by understanding mechanization of knowledge retention within the transformer structure (Dai et al., 2022; Geva et al., 2021; Lv et al., 2024).
@@
We observed that these methods struggle to overcome the challenges of specialized and complex knowledge in the medical domain (Karabacak and Margetis, 2023). As shown in Figure 1, LLMs may exhibit hallucinations and provide answers that do not align with the real world. It lacks specialized knowledge regarding the diseases that can be treated with amoxicillin (Chang et al., 2024; Kung et al., 2023), leading to incorrect responses such as ”stinging”. When dealing with complex knowledge, it is common to employ advanced prompting techniques (e.g., Chain-of-Thought (Lyu et al., 2023)), to facilitate logical reasoning and deduce rational conclusions. However, LLMs may fail to explain the rationale behind the impact of amoxicillin in treating tonsillitis, resulting in an erroneous reasoning chain. Current model editing methods either overlook the storage of complex medical knowledge across different layers in LLMs (Zheng et al., 2023) or introduce substantial modifications to original parameters, which consequently affect the model’s unrelated knowledge and lead to sub-optimal performance (Zhang et al., 2023). This research gap presents a question: Can model editing techniques effectively address challenges when integrating LLMs in the medical domain?
@@
Extensive experiments conducted on two comprehensive benchmarks demonstrate the superior performance of MedLaSA across a range of metrics, showing that it can achieve a higher editing success rate in specialized and complex medical knowledge contexts and avoid impacting irrelevant contexts. To the best of our knowledge, we are the first to propose model editing for enhancing factual knowledge and explanatory abilities in medical LLMs. [^1] https://github.com/quqxui/MedLaSA
@@ 2. Related Work
We present current model editing works following Yao et al. (2023). Memories or Additional Parameters. The methods of with memories or additional parameters typically involve creating explicit memories to store the required knowledge for editing, or adding additional trainable parameters to LLMs for learning new knowledge (Yu et al., 2023; Dong et al., 2022; Hartvigsen et al., 2022; Tang et al., 2022). SERAC (Mitchell et al., 2022b) utilized explicit memory for storing edits and incorporated a scope classifier to understand the editing scope. Given a sample within the editing scope, it utilized a separate model to make edits, ensuring that the original model remains unaffected. IKE (Zheng et al., 2023) designed demonstration formatting and organization strategies, including the copy, update, and retain templates, and retrieved relevant knowledge facts from the editing memories as demonstration inputs to guide the editing process. T-Patcher (Huang et al., 2023) retained all original parameters to preserve overall performance while adding trainable neuron patches to the last Feed-Forward Network (FFN) layer of a Transformer for handling sequential model editing. Despite their success, the above methods lack the exploration of the mechanics of knowledge storage in LLMs, which ultimately leads to poor performance in handling complex medical knowledge. Modifying LLMs’ Parameters. The methods of this category aim to comprehend how knowledge is stored in LLMs and how it can be effectively altered by changing the parameters (De Cao et al., 2021; Geva et al., 2021; Wu et al., 2023). KN (Dai et al., 2022) proposed a knowledge attribution method to identify the neurons associated with specific knowledge without fine-tuning, updating facts, and erasing relations by directly modifying the corresponding parameters in FFN. MEND (Mitchell et al., 2022a) introduced auxiliary hyper-networks to transform the gradient during the fine-tuning process, and trained the hyper-networks to ensure edit success and locality when updating LLMs’ parameters. ROME (Meng et al., 2022a) applied Causal mediation analysis (Pearl, 2022; Vig et al., 2020) to identify decisive neuron activation and modify FFN weights by solving a least squares problem with a linear equality constraint using the Lagrange multiplier. As an extension of ROME (Meng et al., 2022a), MEMIT (Meng et al., 2022b) introduced a multi-layer algorithm to update multiple cases simultaneously. Despite impressive progress made by these methods, they often introduce significant modifications to the original parameters. Consequently, unrelated knowledge is affected, resulting in a noticeable impact on Locality and Fluency, as demonstrated in Section 4 Experiments.
@@
Overall, both of them have their advantages and disadvantages. Neuron localization can achieve interpretability between parameters and knowledge, but sometimes optimizing too much can have negative impacts on unrelated knowledge (Meng et al., 2022a; Geva et al., 2021). On the other hand, adding extra parameters has a smaller impact on unrelated knowledge but lacks exploration of the mechanics of knowledge storage (Huang et al., 2023). Our proposed method, MedLaSA, aims to combine the strengths of both approaches. It identifies the association of knowledge in neurons of different layers, and introduces scalable adapters into the dense layers, enabling precise editing without affecting unrelated knowledge.
@@ 3.1. Prelimimaries
Figure 2. The overview of our MedLaSA. We demonstrate the process of inputting editing, rephrased, and unrelated knowledge.
@@ 3.3. MedLaSA
Our motivation lies in that different knowledge will have different scale sets. Similar contents typically share more consistent scales between them. As a result, it ensures accurate editing of identical knowledge while avoiding impact on unrelated knowledge. For instance, as illustrated in Figure 2, when data that needs editing is inputted, we increase its scale on the $i$-th layer and decrease its impact on the $j$-th layer, ensuring that knowledge updates occur on $i$-th layer. Conversely, when unrelated data is input, the influence of adapter on $j$-th layer is activated and reduces the impact of the adapter on $i$-th layer. Inspired by LoRA (Hu et al., 2021), we incorporate parallel low-rank adapter into dense layers to enable parameter-efficient fine-tuning. Specifically, a pre-trained weight matrix $\boldsymbol{W}_{0}\in\mathbb{R}^{d\times k}$ from attention or MLP module is updated with two trainable matrics $\boldsymbol{B}\in\mathbb{R}^{d\times r}$ and $\boldsymbol{A}\in\mathbb{R}^{r\times k}$. For layer $l$, we have:
@@
The final rank of layer $l$ can be computed by multiplying hyper-parameter $r_{o}$: $r^{l}=\lceil r_{o}\times I^{l}_{r}\rceil$, which ensures the scalability of model editing to different knowledge. By computing scale values $\alpha$ and $r$, we can create a set of layer-wise scaling factors $\{r^{l},\alpha^{l}|l\in[1,L]\}$ for editing, rephrased, and unrelated knowledge, as shown in Figure 2. Besides, a Transformer block typically includes self-attention and MLP modules, whose effects can be separately analyzed with causal tracing. For example, to measure the impact matrix $\boldsymbol{M}_{attn}$ of the attention module, the MLP calculation is cut off and frozen in its corrupted run state, so that it is not affected by the insertion of a clean state. We conducted a thorough analysis to examine the influence of various weights and hyper-parameters on model editing, in Section 4.
@@ 3.4. Medical Model Editing Benchmarks
Explanation / Edit Target: Adiponectin is an abundant adipose-derived protein and enhances insulin sensitivity and lipid oxidation …
@@ 3.4.1. Efficacy and Generality Data Construction
To build the MedCF dataset, we utilize a medical knowledge graph called DRKG (Ioannidis et al., 2020; Xu et al., 2023b). For each triple $(h,r,t)$, we convert $h$ and $r$ into a question and $t$ into an answer, thus creating a QA pair. To evaluate the ability to edit knowledge with unknown prediction results of LLMs, we introduce counterfactual data for model training following ROME (Meng et al., 2022a). We randomly sample a negative entity $t_{*}$ from the knowledge graph and use it as a new answer, attempting to edit the original answer accordingly. For example in Table 1, the original triple “(Primaquine, side effect, Nausea)” is converted to the question: “What side effect is caused by Primaquine?” and the answer “Nausea”. But the editing target is indeed a counterfactual result “Stinging”.
@@
The data of Efficacy for MedCF and Generality for both MedCF and MedFE are generated by querying ChatGPT. The template for querying is shown in Table 2. This prompt template ensures that the generated content adheres to medical terminology while preserving the original meaning of the question. Different from MedCF, all the facts require paraphrasing for the MedFE dataset.
To ensure the credibility and quality of the text generated by ChatGPT, we conducted additional checks. For the MedCF, ChatGPT was employed to convert triplets into questions and rephrase text. As explained in Section 3.4.1, this process involved transforming relation templates, resulting in the need to verify only 34 data points. In the case of the MedFE, questions were already known, and ChatGPT was solely used for rephrasing. In total, there were 4225 data points. We thoroughly examined all the generated data and took measures to ensure that the rephrased sentences maintained their original meaning.
@@
Table 2. Templates for querying ChatGPT (gpt-3.5-turbo-0613) to generate the Efficacy and Generality data. $h$, $r$, and $t$ denote the head entity, relation, and tail entity, respectively.
@@
The QA pair of Locality ($x_{i}$,$y_{i}$) aims to assess the impact of model editing on unrelated knowledge. Previous benchmarks either employed out-of-distribution data (e.g., zsRE (Mitchell et al., 2022a)) or solely relied on data with the same ground truth (e.g., CounterFact (Meng et al., 2022a)). Nevertheless, we argue that a comprehensive evaluation of Locality is necessary to prevent the inadvertent modification of irrelevant knowledge and ensure high reliability of the medical domain. As we explained in Section 1, the post-edited model should be evaluated based on the following challenges: Target Distribution (TD), Entity Mapping (EM), Structural Similarity (SS), Textual Similarity (TS), and Consistent Topic (CT). The issues that these evaluation metrics aim to address are mentioned in Section 1. To fulfill the requirements, we gathered relevant QA pairs for evaluation. The original triple of Efficacy QA pair is denoted as $(h,r,t)$. We employ the graph structure of KG for searching and building Locality sub-metrics. For Locality-TD, we sample new triples $(h_{td},r_{td},t)$ by fixing $t$ to investigate the shift of the ground-truth token. For Locality-EM, we sample new triples $(h,r_{e}m,t_{e}m)$ by fixing $h$ to create mappings between head and tail entities. For Locality-SS, we utilize the knowledge graph embedding method, RotatE (Sun et al., 2018), to learn the entity and relation embeddings from the graph structure. For Locality-TD, we use BioBERT (Lee et al., 2020) to convert $(h,r,t)$ triples into text embeddings, and then obtain similar triples $(h_{s},r_{s},t_{s})$ based on Structural Similarity and Textual Similarity, respectively. For Locality-CT, we utilize the source data from MedMCQA, which includes topic tags for each question, annotated by experts. These new triples are also converted into QA pairs, which enable a thorough analysis of the impacts of model editing techniques on unrelated knowledge in medical domain. It should be noted that the calculation for all sub-metrics of Locality is the same, as listed in Section 3.1.
We provide knowledge type distribution of MedCF and MedFE in Figure 3, respectively. The knowledge types in MedCF are categorized based on their relations, while the types in MedFE are categorized according to medical subjects. Data splits of MedCF and MedFE are shown in Table 3. We have attempted to ensure that different categories have similar proportions, but due to data limitations, some types have not been collected as extensively as others. More details about datasets can be found in our repository.